The table below shows the national income statistics for a country.

Items / RM (Million)

Import / RM 6700

Export / RM 8500

Subsidies / RM 480

Consumption / RM 3400

Indirect taxes/ RM 860

Depreciation/ RM 520

Government Expenditure/ RM 4700

Investment / RM 6900

Factor income received abroad/ RM 2010

Corporate tax / RM 620

Factor income paid to abroad/ RM 1800

From the data, calculate:

a) Gross Domestic Product at market price

b) Gross National Product at market price

c) Gross National Product at factor cost

d) Gross Domestic Product at factor cost

e) National Income


1
Expert's answer
2022-01-31T07:38:45-0500

Solution:

a.). Gross Domestic Product at market price = Consumption + Government Expenditure + Investment + Corporate tax + Net exports = 3,400 + 4,700 + 6,900 + 620 + (8,500 – 6,700) = RM 17,420


b.). Gross National Product at market price = GDP at market price + Factor income received abroad = 17,420 + 2,010 = RM 19,430

 

c.). Gross National Product at factor cost = GDP at market price – Indirect taxes + Factor income paid to abroad = 17,420 – 860 + 1,800 = RM 18,360

 

d.). Gross Domestic Product at factor cost = GDP at market price – Indirect taxes = 17,420 – 860 = RM 16,560

 

e.). National income = C + G + I + X – M = 3,400 + 4,700 + 6,900 + (8,500 – 6,700) = RM 16,800